Use the elimination method to find all solutions of the system of equations. (Round each answer to two decimal places.)
x^2 - y^2 = 25
2x^2 - y^2 = x + 55

Answer by checkley77(12844)

x^2 - y^2 = 25
-(2x^2 - y^2 = x + 55) subtract
-------------------------------
-x^2=-x+25-55
-x^2+x+30=0
x^2-x-30=0
(x-6)(x+5)=0
x-6=0
x=6 ans. (6,+-3)
x=5=0
x=-5 ans.
6^2-y^2=25
36-y^2=25
-y^2=25-36
-y^2=-9
y^2=9
y=+-3 ans.
2*-5^2-y^2=-5+55
2*25-y^2=50
50-y^2=50
-y^2=0 ans. (-5,0)